OpenStudy (anonymous):

I'm supposed to find y'' in terms of x and y given x^5+y^5=2^5 I know how to do the first derivative, but I'm struggling with the second...

OpenStudy (btaylor):

what do you have for the first derivative?

OpenStudy (anonymous):

5x^4+5y^4*y'=0

OpenStudy (anonymous):

for the second my guess is 20x^3+20y^3*y'+5y^4*y''=0...

OpenStudy (btaylor):

That looks right, but you need to solve for y' and plug it into that equation. Then solve that giant mess for y''.

OpenStudy (anonymous):

so y' will be -x^4/y^4..

OpenStudy (btaylor):

yep.

OpenStudy (anonymous):

i get (4x^4/y-4x^3)/y^4 ?

OpenStudy (anonymous):

that's y''= (4x^4/y-4x^3)/y^4

OpenStudy (anonymous):

I'm still doing something wrong...

OpenStudy (btaylor):

i think that's right...

OpenStudy (anonymous):

wait, when I take the derivative of y' it doesn't equal y'', it's more than that...do you know what it is @BTaylor?

zepdrix (zepdrix):

No it's not more than that.\[\Large\bf\sf \frac{d}{dx}\frac{dy}{dx}\quad=\quad \frac{d^2y}{dx^2}\quad=\quad y''\]

OpenStudy (isaiah.feynman):

@josh55 are you clarified?